Consider the following time series data.

t 1 2 3 4 5
Yt 6 12 9 14 16

a. Which of the following is a correct time series plot for this data?

time series plot #1

time series plot #2

time series plot #3

1.
2.
3.

What type of pattern exists in the data?

linear trend

curvilinear trend

horizontal

b. Develop the linear trend equation for this time series (to 1 decimal).

Tt = ___ + ____t

c. What is the forecast for t=6 (to 1 decimal)?
Answer #1

a
)

2345

type of pattern exists in the data : linear trend

b)

1 29.16 12 0.36 3 (x-x) 4.0000 1.0000 0.0000 10000 4.0000 10.0000 SSX 14 16 5.76 6.76 21.16 63.20 SSY 57 Total Mean 3.000 11.

from above

Tt =4.8+2.2t

c)

predicted val=4.8+6*2.2= 18.000